# Center vortices in the novel phase of staggered fermions

### v1

@@ Abstract
The geometry of center vortices is studied in the novel lattice-artefact phase that appears with staggered fermions to elucidate any insight provided by the center-vortex degrees of freedom. For various numbers of fermion flavors, the single-site shift symmetry of the staggered-fermion action is broken in a finite region of the $(\beta,m)$ phase space. Simulations are performed with six degenerate fermion flavors and a range of $\beta$ values that span the phase boundary. Center vortices are demonstrated to capture the broken shift symmetry that manifests in the unphysical phase. This persists at the level of each individual plaquette orientation, where it is revealed only the plaquettes that span the broken dimension are affected. Several bulk center-vortex quantities, including the vortex and branching point densities, are considered to highlight other aspects of vortex geometry sensitive to the unphysical phase. A slight preference for the plaquettes affected by the broken shift symmetry to be pierced by a vortex is observed. This translates also to a greater branching point density in three-dimensional slices that span the broken dimension. Combined, these findings provide a novel characterization of the unphysical phase in terms of the fundamental center degrees of freedom.
@@ I Introduction
The possibility of an unphysical lattice-artefact phase in the $(\beta,m)$ phase diagram was noted in Refs. [1, 2]. Such a phase should naturally be avoided in taking the continuum limit. Concrete simulation results demonstrating the presence of the unphysical phase were reported in Refs. [3, 4, 5] with various flavor content, all with gauge group $\mathrm{SU}(3)$ and some amount of Dirac operator improvement in the form of smearing. Taking all quark flavors degenerate, the unphysical phase corresponds to a finite region in the $(\beta,m)$ plane as illustrated in Fig. 1 for various flavor numbers. The unphysical phase has also been studied within the framework of chiral perturbation theory in Ref. [6]. Interestingly, there is currently no convincing evidence for the presence of the unphysical phase with unimproved, unsmeared staggered fermions.
Figure 1: Adapted from Fig. 2 in Ref. [5]. Approximate locations of the broken-shift-symmetry phase boundaries in the $(\beta,m)$ plane for various numbers of fermion flavors $N_{f}$. The phase is located under the curves and extending down to a bare quark mass $m=0$. The thickness of the curves reflects an uncertainty in the precise boundary locations. A transition does exist for $N_{f}=3$, but cannot be properly resolved at the shown scale.
@@
while it is intact in the physical phase connected to the continuum limit. Here $\chi(x)$ is the staggered fermion field on the lattice with signs $\xi_{\mu}(x)=(-1)^{\sum_{\nu>\mu}x_{\nu}}$, and $U_{\mu}(x)$ are the usual link variables.
In this work, we examine this unphysical phase from the viewpoint of center vortices [8, 9, 10, 11]. We aim to determine if center vortices can capture this physics and if so, learn the insight provided through the relatively simple description of nonperturbative phenomena in QCD.
Extensive studies have demonstrated that the center-vortex degrees of freedom underpin many nonperturbative phenomena of lattice gauge theory, including confinement and chiral symmetry breaking [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. In recent work, the dependence of center vortices on the number of flavors was analysed in relation to the conformal window [45], where care had to be taken to avoid the unphysical phase. It is thus of interest to investigate to what extent the vortex degrees of freedom capture the broken-shift-symmetry phase. If successful, significant insight into the unphysical phase may be obtained.
@@ II Center vortices
To investigate center vortices in the unphysical phase, six degenerate flavors of staggered fermions are employed with a bare mass $m=0.015$ and $\beta$ values ranging from 2.52.52.52.52.92.92.92.9. These parameters are seen from Fig. 1 to span the boundary of the broken-shift-symmetry phase for $N_{f}=6$. Symmetric $16^{4}$ lattice volumes are utilized over the full $\beta$ range, with $24^{4}$ volumes also considered for a subset of $\beta$ values to demonstrate that our results are not afflicted by finite volume effects.
@@ III Unphysical phase
where $P_{\mu\nu}(x)=U_{\mu}(x)\,U_{\nu}(x+\hat{\mu})\,U_{\mu}^{\dagger}(x+\hat{\nu})\,U_{\nu}^{\dagger}(x)$, and for each $\mu$ the average in Eq. (6) is taken only over even lattice sites in the $\mu$ direction. The quantity $P(x)$ is a sum over all plaquette orientations at $x$. In the physical phase, we expect $\Delta_{\mu}P=0$ for each $\mu$, with the understanding that there is no preferred sign for the difference $P(x)-P(x+\hat{\mu})$ between adjacent plaquettes. In the unphysical phase, however, $\Delta_{\mu}P$ develops a nonzero value along one or more dimensions.
@@
Turning to the vortex-only fields (Fig. 3), we observe similar qualitative behavior in $\left(\Delta_{\mu}P\right)^{2}$. Namely, it starts off small at large $\beta$ and increases towards the left. This immediately indicates that the center-vortex degrees of freedom are sufficient to capture the broken shift symmetry. It is interesting to note that the value of $\left(\Delta_{\mu}P\right)^{2}$ is approximately an order of magnitude larger on the vortex-only fields compared to the untouched configurations. This has a simple explanation. In the vortex fields, the difference between adjacent plaquettes is only nonzero when one plaquette is pierced by a vortex while the other is trivial. Owing to the discrete nature of the center, the smallest possible such difference is therefore ±$\pm 1.5$. This is in contrast to the untouched configurations where the plaquette takes a continuous spectrum, and as such the difference $P(x)-P(x+\hat{\mu})$ tends to be smaller.
Having shown at a high level that center vortices are sensitive to the unphysical phase, we now wish to consider the broken symmetry in greater detail. It was mentioned that $\Delta_{\mu}P$ can be nonzero in one or more dimensions at a time, and additionally this dimension can vary as a function of simulation time (i.e. on a configuration-to-configuration basis). To examine this, we select one $24^{4}$ ensemble in the unphysical phase at $\beta=2.675$, and calculate $\Delta_{\mu}P$ for each $\mu$ under a moving average of thirty configurations. This window size is chosen to be sufficiently large to produce a smooth curve as the window is shifted, while also being sufficiently small to demonstrate the variation in $\Delta_{\mu}P$ with simulation time. The resulting evolution is presented for the untouched configurations in Fig. 5, and again for the vortex-only fields in Fig. 5.
Figure 4: A moving average of $\Delta_{\mu}P$ over thirty configurations for each dimension $\mu$, calculated on the untouched fields. This demonstrates how the shift symmetry can be broken along multiple dimensions simultaneously, and additionally that the strength of the symmetry breaking can vary with simulation time.
@@
Initially, the symmetry breaking is most pronounced in the temporal dimension $(\mu=t)$, for which the difference $\Delta_{\mu}P$ is negative. This is seen to weaken as the thirty-configuration window is shifted, while simultaneously becoming stronger in the $y$ dimension where the difference is positive. This illustrates how the symmetry breaking can move between the dimensions, and is not locked into a single sign. Furthermore, there is a mild breaking observed in the $x$ and $z$ dimensions over a majority of the analyzed range. Nearly identical behavior is exhibited on both the untouched and vortex-only configurations, though the value of $\Delta_{\mu}P$ is again considerably larger on the latter.
@@
Here, we discover that it is in fact only the plaquettes that include the dimension under consideration for which the shift symmetry is broken. The individual order parameters for $xy$, $xt$ and $yt$ plaquettes, none of which include the $z$ dimension, are consistent with zero over the entire considered range. On the other hand, the equivalent computations for $xz$, $yz$ and $zt$ plaquettes, all of which cover the $z$ dimension, reveal persistently nonzero values of $\Delta_{z}P_{\mu\nu}$, indicating breaking of the shift symmetry. The order parameter obtained from averaging over all plaquette orientation unsurprisingly lies halfway between the two groups.
@@
with the ratio defined to be $>1$ when $\Delta_{\mu}P$ is positive.
@@
Similar qualitative features are exhibited as in Fig. 7. Namely, for the plaquette orientations that do not span the broken $z$ dimension ($xy$, $xt$, $yt$) the ratio is approximately unity. This indicates there is no preference for even or odd plaquettes, and that the shift symmetry is in tact. For the other plaquettes that do contain the $z$ dimension, the ratio is notably greater than 1, a clear signal of the asymmetry.
However, the primary purpose of computing these ratios is to quantify the extent of the even-odd asymmetry. We see from Fig. 8 that the maximum value attained within one window of configurations is $\simeq 1.1$, signifying a 10% asymmetry. That is to say, for every 10 even plaquettes pierced by a vortex, there are 11 odd plaquettes pierced. Unfortunately, it does not seem likely that such a discrepancy will be identifiable in visualizations. The maximum ratio on any single configuration was $\simeq 1.15$, which although slightly larger still seems insufficient. Recalling that the ensemble used to produce these plots was chosen due to its maximal symmetry breaking, we therefore concede that the broken shift symmetry is not likely to be gleaned from visualizations.
@@ IV.1 Vortex density
Remarkably, there is a suggestion of a discrepancy between the different plaquette orientations. That is, the $xz$, $yz$ and $zt$ densities appear to sit above the $xy$, $xt$ and $yt$ densities over the entire range. Recalling it is the $z$ dimension along which the symmetry is broken, this is precisely the manner in which a difference might appear. To aid in verifying whether this is genuine, we reproduce the plot in Fig. 11 averaging over the three plaquette orientations within each group. This smooths out some of the noise present in Fig. 11, and reveals clearly a greater density of pierced plaquettes containing the broken $z$ dimension. Furthermore, the split persists over all 100 configurations; this is the purpose of continuing to utilize a moving average.
@@ IV.2 Branching point density
Figure 12: Schematic of a monopole vertex (left) versus a branching point (right). The monopole shows the directed flow of $m=+1$ center charge. Reversal of the left-hand arrow indicates the flow of $m=-1$ charge, as seen on the right. Due to periodicity in the center charge, $m=-1$ is equivalent to $m=+2$. Thus, the right-hand diagram depicts the branching of center charge.
@@
We wish to determine whether a similar divergence as in the densities of plaquettes pierced exists for branching points. In this case, the necessary division is into the three-dimensional slices that contain the broken dimension $\mu$ and the one slice that does not, obtained from slicing through the broken dimension itself (i.e. holding the $\mu$ coordinate fixed). The branching point densities for these two cases are presented as a moving average in Fig. 14. We use the terminology “$\mu$ slice” to refer to the three-dimensional slice obtaining by fixing the $\mu$ coordinate.
Figure 14: The branching point densities for fixed-$z$ and combined fixed-$x,y,t$ slices under a moving average for $\beta=2.55$. We find that the branching point density in three-dimensional slices that span the broken $z$ dimension is persistently larger, in line with the greater density of $xz$, $yz$ and $zt$ plaquettes pierced.
@@ V Conclusion
In this work, we have studied center vortices in the lattice-artefact phase of staggered fermions, in which the single-site shift symmetry of the staggered-fermion action is broken in a finite region of the $(\beta,m)$ phase space. Simulations with six degenerate fermion flavors at $\beta$ values ranging between 2.52.52.52.52.92.92.92.9 and lattice volumes of $16^{4}$ and $24^{4}$ were employed for the study.
Center vortices were verified to capture the broken shift symmetry manifest in the unphysical phase, as quantified by the order parameter $\Delta_{\mu}P$ in Eq. (6). This indicated there is a preference for either even or odd plaquettes to be pierced by a vortex in the unphysical phase. The behavior of $\Delta_{\mu}P$ along each dimension was examined, demonstrating that the shift symmetry can be broken along multiple dimensions simultaneously and moreover that the degree of symmetry breaking along a given dimension can change with simulation time.
The order parameter was further decomposed down to the individual plaquette level, revealing for the first time that only the plaquette orientations spanning the broken dimension are affected. This was established to hold for both the untouched and vortex fields, certifying that center vortices capture the symmetry breaking at the most granular level. Ratios of the number of odd to even plaquettes pierced by a vortex revealed a best-case asymmetry of $\simeq 10\%$ on our ensembles, making direct observation through visualizations difficult.
@@ Acknowledgements.
This work was supported with supercomputing resources provided by the Phoenix HPC service at the University of Adelaide. This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. This research was supported by the Australian Research Council through Grant No. DP210103706. D. N. was supported by the NKFIH under grants Excellence No. TKP2021-NKTA-64, NKKP Excellence No. 151482 and K-147396. D. N. acknowledges support from the George Southgate Fellowship and thanks the CSSM and Department of Physics at the University of Adelaide for their kind hospitality during his visit where parts of this work were completed.

### v2

@@ Abstract
The geometry of center vortices is studied in the novel lattice-artefact phase that appears with staggered fermions to elucidate any insight provided by the center-vortex degrees of freedom. For various numbers of fermion flavors, the single-site shift symmetry of the staggered-fermion action is broken in a finite region of the $(\beta,m)$ phase space. Simulations are performed with six degenerate fermion flavors and a range of $\beta$ values that span the phase boundary. Center vortices are demonstrated to capture the broken shift symmetry that manifests in the unphysical phase. This persists at the level of each individual plaquette orientation, where it is revealed that only the plaquettes that span the broken dimension are affected. Several bulk center-vortex quantities, including the vortex and branching point densities, are considered to highlight other aspects of vortex geometry sensitive to the unphysical phase. A slight preference for the plaquettes affected by the broken shift symmetry to be pierced by a vortex is observed. This translates also to a greater branching point density in three-dimensional slices that span the broken dimension. Combined, these findings provide a novel characterization of the unphysical phase in terms of the fundamental center degrees of freedom.
@@ I Introduction
The possibility of an unphysical lattice-artefact phase in the $(\beta,m)$ phase diagram was noted in Refs. [1, 2]. Such a phase should naturally be avoided in taking the continuum limit. Concrete simulation results demonstrating the presence of the unphysical phase were reported in Refs. [3, 4, 5] with various flavor content, all with gauge group $\mathrm{SU}(3)$ and some amount of Dirac operator improvement in the form of smearing. Taking all quark flavors as degenerate, the unphysical phase corresponds to a finite region in the $(\beta,m)$ plane, as illustrated in Fig. 1 for various flavor numbers. The unphysical phase has also been studied within the framework of chiral perturbation theory in Ref. [6]. Interestingly, there is currently no convincing evidence for the presence of the unphysical phase with unimproved, unsmeared staggered fermions.
Figure 1: Adapted from Fig. 2 in Ref. [5]. Approximate locations of the broken-shift-symmetry phase boundaries in the $(\beta,m)$ plane for various numbers of fermion flavors $N_{f}$. The phase is located under the curves and extends down to a bare quark mass $m=0$. The thickness of the curves reflects an uncertainty in the precise boundary locations. A transition does exist for $N_{f}=3$, but cannot be properly resolved at the shown scale.
@@
while it is intact in the physical phase connected to the continuum limit. Here, $\chi(x)$ is the staggered fermion field on the lattice with signs $\xi_{\mu}(x)=(-1)^{\sum_{\nu>\mu}x_{\nu}}$, and $U_{\mu}(x)$ are the usual link variables.
In this work, we examine this unphysical phase from the viewpoint of center vortices [8, 9, 10, 11]. We aim to determine if center vortices can capture this physics and, if so, learn the insight provided through the relatively simple description of nonperturbative phenomena in QCD.
Extensive studies have demonstrated that the center-vortex degrees of freedom underpin many nonperturbative phenomena of lattice gauge theory, including confinement and chiral symmetry breaking [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. In recent work, the dependence of center vortices on the number of flavors was analyzed in relation to the conformal window [45], where care had to be taken to avoid the unphysical phase. It is thus of interest to investigate to what extent the vortex degrees of freedom capture the broken-shift-symmetry phase. If successful, significant insight into the unphysical phase may be obtained.
@@ II Center vortices
To investigate center vortices in the unphysical phase, six degenerate flavors of staggered fermions are employed with a bare mass $m=0.015$ and $\beta$ values ranging from 2.52.52.92.9. These parameters are seen from Fig. 1 to span the boundary of the broken-shift-symmetry phase for $N_{f}=6$. Symmetric $16^{4}$ lattice volumes are utilized over the full $\beta$ range, with $24^{4}$ volumes also considered for a subset of $\beta$ values to demonstrate that our results are not afflicted by finite volume effects.
@@ III Unphysical phase
[…]
@@
[…]
Having shown at a high level that center vortices are sensitive to the unphysical phase, we now wish to consider the broken symmetry in greater detail. It was mentioned that $\Delta_{\mu}P$ can be nonzero in one or more dimensions at a time, and additionally this dimension can vary as a function of simulation time (i.e. on a configuration-to-configuration basis).
To examine this, we select one $24^{4}$ ensemble in the unphysical phase at $\beta=2.675$, and calculate $\Delta_{\mu}P$ for each $\mu$ under a moving average of 30 configurations in the HMC trajectory. This window size is chosen to be sufficiently large to produce a smooth curve as the window is shifted, while also being sufficiently small to demonstrate the variation in $\Delta_{\mu}P$ with simulation time. To be precise, the value associated with the $i$th window of 30 configurations is calculated by averaging over configurations $i$ through $i+29$ in simulation time, with ten HMC trajectories between each sampled configuration. The resulting evolution is presented for the untouched configurations in Fig. 5, and again for the vortex-only fields in Fig. 5.
Figure 4: A moving average of $\Delta_{\mu}P$ over a window of 30 configurations in simulation time for each dimension $\mu$, calculated on the untouched fields. This demonstrates how the shift symmetry can be broken along multiple dimensions simultaneously, and additionally that the strength of the symmetry breaking can vary with simulation time.
@@
Initially, the symmetry breaking is most pronounced in the temporal dimension $(\mu=t)$, for which the difference $\Delta_{\mu}P$ is negative. This is seen to weaken as the 30-configuration window is shifted, while simultaneously becoming stronger in the $y$ dimension where the difference is positive. This illustrates how the symmetry breaking can move between the dimensions, and is not locked into a single sign. Furthermore, there is a mild breaking observed in the $x$ and $z$ dimensions over a majority of the analyzed range. Nearly identical behavior is exhibited on both the untouched and vortex-only configurations, though the value of $\Delta_{\mu}P$ is again considerably larger on the latter.
@@
Here, we discover that it is in fact only the plaquettes that include the dimension under consideration for which the shift symmetry is broken. The individual order parameters for $xy$, xtxt, and $yt$ plaquettes, none of which include the $z$ dimension, are consistent with zero over the entire considered range. On the other hand, the equivalent computations for $xz$, yzyz and $zt$ plaquettes, all of which cover the $z$ dimension, reveal persistently nonzero values of $\Delta_{z}P_{\mu\nu}$, indicating breaking of the shift symmetry. The order parameter obtained from averaging over all plaquette orientation unsurprisingly lies halfway between the two groups.
@@
with the ratio defined to be >1>1 when $\Delta_{\mu}P$ is positive.
@@
Similar qualitative features are exhibited as in Fig. 7. Namely, for the plaquette orientations that do not span the broken $z$ dimension ($xy$, xtxt, ytyt), the ratio is approximately unity. This indicates there is no preference for even or odd plaquettes, and that the shift symmetry is intact. For the other plaquettes that do contain the $z$ dimension, the ratio is notably greater than 1, a clear signal of the asymmetry.
However, the primary purpose of computing these ratios is to quantify the extent of the even-odd asymmetry. We see from Fig. 8 that the maximum value attained within one window of configurations is $\simeq 1.1$, signifying a 10% asymmetry. That is to say, for every 10 even plaquettes pierced by a vortex, there are 11 odd plaquettes pierced. Unfortunately, it does not seem likely that such a discrepancy will be identifiable in visualizations. The maximum ratio on any single configuration was $\simeq 1.15$, which, although slightly larger, still seems insufficient. Recalling that the ensemble used to produce these plots was chosen due to its maximal symmetry breaking, we therefore concede that the broken shift symmetry is not likely to be gleaned from visualizations.
@@ IV.1 Vortex density
Remarkably, there is a suggestion of a discrepancy between the different plaquette orientations. That is, the $xz$, yzyz, and $zt$ densities appear to sit above the $xy$, xtxt and $yt$ densities over the entire range. Recalling it is the $z$ dimension along which the symmetry is broken, this is precisely the manner in which a difference might appear. To aid in verifying whether this is genuine, we reproduce the plot in Fig. 11 averaging over the three plaquette orientations within each group. This smooths out some of the noise present in Fig. 11, and reveals clearly a greater density of pierced plaquettes containing the broken $z$ dimension. Furthermore, the split persists over all 100 configurations; this is the purpose of continuing to utilize a moving average.
@@ IV.2 Branching point density
Figure 12: Schematic of a monopole vertex (left) versus a branching point (right). The monopole shows the directed flow of $m=+1$ center charge. Reversal of the left-hand arrow indicates the flow of m=1m=-1 charge, as seen on the right. Due to periodicity in the center charge, m=1m=-1 is equivalent to $m=+2$. Thus, the right-hand diagram depicts the branching of center charge.
@@
We wish to determine whether a similar divergence as in the densities of plaquettes pierced exists for branching points. In this case, the necessary division is into the three-dimensional slices that contain the broken dimension $\mu$ and the one slice that does not, obtained from slicing through the broken dimension itself (i.e. holding the $\mu$ coordinate fixed). The branching point densities for these two cases are presented as a moving average in Fig. 14. We use the terminology $\mu$ slice” for the three-dimensional slice obtaining by fixing the $\mu$ coordinate.
Figure 14: The branching point densities for fixed-$z$ and combined fixed-$x,y,t$ slices under a moving average for $\beta=2.55$. We find that the branching point density in three-dimensional slices that span the broken $z$ dimension is persistently larger, in line with the greater density of $xz$, yzyz, and ztzt plaquettes pierced.
@@ V Conclusion
In this work, we have studied center vortices in the lattice-artefact phase of staggered fermions, in which the single-site shift symmetry of the staggered-fermion action is broken in a finite region of the $(\beta,m)$ phase space. Simulations with six degenerate fermion flavors at $\beta$ values ranging between 2.52.52.92.9, and lattice volumes of $16^{4}$ and $24^{4}$ were employed for the study.
Center vortices were verified to capture the broken shift symmetry manifest in the unphysical phase, as quantified by the order parameter $\Delta_{\mu}P$ in Eq. (6). This indicated there is a preference for either even or odd plaquettes to be pierced by a vortex in the unphysical phase. The behavior of $\Delta_{\mu}P$ along each dimension was examined, demonstrating that the shift symmetry can be broken along multiple dimensions simultaneously and, moreover, that the degree of symmetry breaking along a given dimension can change with simulation time.
The order parameter was further decomposed down to the individual plaquette level, revealing for the first time that only the plaquette orientations spanning the broken dimension are affected. This was established to hold for both the untouched and vortex fields, certifying that center vortices capture the symmetry breaking at the most granular level. Ratios of the number of odd to even plaquettes pierced by a vortex revealed a best-case asymmetry of $\simeq 10\%$ in our ensembles, making direct observation through visualizations difficult.
@@ Acknowledgements.
D. N. acknowledges support from the George Southgate Fellowship and thanks the CSSM and Department of Physics at the University of Adelaide for their kind hospitality during his visit where parts of this work were completed. This work was supported with supercomputing resources provided by the Phoenix HPC service at the University of Adelaide. This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. This research was supported by the Australian Research Council through Grant No. DP210103706. D. N. was supported by the NKFIH under grants Excellence No. TKP2021-NKTA-64, NKKP Excellence No. 151482 and K-147396.